Solve the equation 4 times cosine of theta plus 1 = 2 times cosine of theta for theta greater than or equal to 0 but less than 2 pi..
Answer by stanbon(75887)

Solve the equation 4 times cosine of theta plus 1 = 2 times cosine of theta for theta greater than or equal to 0 but less than 2 pi..
------
4cos(t)+1 = 2cos(t) for 0<= t <2pi
-----------------------
2cos(t) = -1
-----
cos(t) = -1/2
-----
t = 180-60 or 180+60
t = 120 degrees or 240 degrees
t = (2/3)pi or (4/3)pi radians
-----------------------------------
Cheers,
Stan H.
========================